# Orchestrating Rewards in the Era of Intelligence-Driven Commerce

### v1

@@ Abstract
Keywords: Loyalty programs; mechanism design; platform economics; automated market makers; pricing mechanisms; coalition programs
@@ 1 Introduction
Coalition loyalty programs emerged as the dominant attempt to implement this vision, with multiple brands partnering through central operators who manage shared currencies that customers earn and redeem across partner networks. The theoretical advantages appeared compelling, rooted in network economics, economies of scale, and enhanced customer value propositions. The empirical reality, however, proved starkly different. Industry analysis reveals that approximately 60% of coalition programs fail within 10 years of launch, with high-profile examples including American Express’s Plenti, which shut down after three years despite over \$100 million in investment (Bond Brand Loyalty, 2020; Peterson, 2018).
This widespread failure pattern raises a fundamental question that motivates our research. Do coalitions face merely operational challenges that better management could resolve, or do they suffer from structural problems inherent to their architectural design? Our systematic analysis reveals that coalition failures stem from fundamental architectural limitations in centralized operator models rather than operational deficiencies. These structural problems grow increasingly critical as commerce transforms, with artificial intelligence systems now mediating relationships between consumers and businesses through shopping assistants, browser extensions, and autonomous purchasing agents (Russell and Norvig, 2020). When intelligent agents require standardized, machine-readable interfaces for effective operation, the architectural limitations of traditional coalition models become absolute barriers rather than mere inefficiencies.
@@ 2.2.1 Closed Systems
However, closed systems face inherent limitations that restrict their value proposition. Customers experience fragmentation, managing multiple programs with separate accounts, balances, and rules, creating cognitive burden and reducing engagement (Nunes and Drèze, 2006). Limited network effects emerge because value remains constrained to single-brand touchpoints, preventing the accumulation synergies that multi-brand systems could theoretically provide. Smaller businesses struggle with economies of scale, finding the infrastructure investments and operational costs prohibitive relative to their customer base and transaction volumes. Breakage (unredeemed points) may be high when customers lack sufficient engagement with single brands to accumulate meaningful balances, representing deadweight loss in the system. Most fundamentally, closed systems misapprehend the reality that users are inherently social beings. The same individual who flies an airline also dines at restaurants, shops at malls, stays in hotels, and attends sporting events. While such systems may suffice in today’s landscape, they face inevitable obsolescence as thousands of businesses compete for increasingly scarce user attention. In this fragmented environment, requiring users to independently manage rewards across disparate platforms becomes untenable. Moreover, in an emerging AI-driven agentic paradigm capable of computing complex user preferences instantaneously, closed systems dramatically underutilize optimization capabilities. Strategies that optimize across composite user goals, treating users as the social entities they are rather than isolated economic actors, will demonstrably outperform siloed approaches that fail to reflect the interconnected nature of human behavior.
@@ 2.3.1 Empirical Evidence from Coalition Failures
Air Miles Canada, one of the longest-running coalition programs, grew to over 10 million member households and more than 100 brand partners over two decades, achieving scale that should have generated the network effects and economies central to coalition value propositions. However, the program faced persistent profitability challenges that led to a critical 2016 decision where announcing that points would begin expiring triggered massive customer backlash, government scrutiny, and intense negative media coverage (CBC News, 2016). The coalition operator was ultimately forced to reverse the policy, but the incident severely damaged program trust and exposed fundamental tensions between coalition economics and customer satisfaction. While Air Miles survived this crisis, the experience revealed that even large-scale, established coalitions struggle with economic sustainability.
Nectar in the United Kingdom launched in 2002 with prominent anchor partners including Sainsbury’s, BP, and Barclaycard, rapidly accumulating 19 million members and achieving high recognition and participation rates in its early years (Uncles et al., 2003). The coalition appeared to validate the open systems thesis, demonstrating that cross-brand reward ecosystems could achieve massive scale. However, the program gradually lost major partners and experienced declining engagement as the economic reality failed to match theoretical projections. By 2019, primary partner Sainsbury’s announced plans to replace Nectar with a proprietary loyalty program (Wood, 2019), signaling that even successful coalitions struggle to maintain partner commitment long-term. The gradual erosion of the partner network suggests that coalition participation does not deliver sustainable value propositions for major brands.
@@
Plenti in the United States represents perhaps the most dramatic coalition failure. Backed by American Express and featuring high-profile partners including AT&T, Exxon, Macy’s, and Rite Aid, Plenti launched in 2015 with over \$100 million in marketing investment and ambitious projections to establish the first major multi-brand loyalty program in the U.S. market (Peterson, 2018). However, within only three years, American Express announced the program’s shutdown, citing inability to achieve projected growth and engagement targets. The rapid failure despite enormous resources, sophisticated operator expertise, and prominent partners demonstrated conclusively that coalition challenges extend beyond operational execution to fundamental structural issues that cannot be resolved through better management or increased investment.
@@ 2.5 Trustless Perpetual Systems: Architectural Foundations for Alternative Coordination
Automated Market Makers represent a specific smart contract application particularly relevant to reward systems. AMMs use algorithmic formulas to determine exchange rates between assets based on supply and demand reflected in liquidity pools (Adams et al., 2020). Participants provide liquidity by depositing assets, earning fees in exchange, while others execute trades at algorithmically determined rates. This mechanism enables decentralized exchange without operators setting prices, matching counterparties, or maintaining order books.
@@ 4.2.2 Dynamic Adjustment Factors
Flow Balance and Liability Management ($\phi_{A}$). When brands consistently accept more rewards than their customers redeem elsewhere, they accumulate liabilities that create financial risk and potential redemption obligations exceeding available resources (Chun and Iancu, 2019). The flow balance factor $\phi_{A}=(\text{Outflow}_{A}-\text{Inflow}_{A})/(\text{Outflow}_{A}+\text{Inflow}_{A})$ captures this dynamic, providing pricing signals that encourage reciprocal flows and discourage persistent imbalances. Brands with excessive outflows face premiums discouraging further redemptions elsewhere, while brands with excessive inflows receive discounts encouraging their customers to redeem at other brands.
Transaction Size and Systemic Risk ($\mu$). Large transactions relative to available system capacity create systemic risk affecting subsequent exchange availability and pricing stability (Chun and Iancu, 2019). A single transaction exhausting significant capacity impacts other users’ ability to execute exchanges. The transaction size factor $\mu=V_{\text{transaction}}/C_{\text{system}}$, where $V_{\text{transaction}}$ represents the transaction’s value and $C_{\text{system}}$ represents total system capacity, enables dynamic adjustment based on relative transaction magnitude. Large transactions face premiums reflecting their disproportionate resource consumption.
@@ 4.7.6 Withdrawal Mechanisms and Safeguards
5 Pricing Mechanism: Validation and Operationalization
5.1 Validation Against Loyalty Program Success Criteria
The comprehensive customer pricing model developed in Section 4.2 incorporates multiple adjustment factors to account for the complex economic dynamics of cross-brand reward exchange. This model can be expressed as:
Expanding with the identified factors yields:
$\quad\times\left[1+\beta_{\text{quality}}\times(1-\rho_{B})\right]\times[1+\Omega]+\varepsilon$
where the eight factors address transaction risk, aggregate flows, brand synergies, competitive effects, market conditions, temporal variations, service quality, and customer attributes respectively.
The fundamental validation question we must address is whether this pricing mechanism can satisfy the dual requirements for loyalty program success: maintaining customer satisfaction through perceived fairness and value while ensuring brand participation through sustainable economics. Extensive research demonstrates that pricing adjustments aligned with customer behavior and brand economics significantly enhance program performance. Firms in coalition loyalty programs achieve profitability when consumers earn rewards broadly across partners and when reward managers can account for individual firm pricing in setting values (Allaway et al., 2006). Studies on integrated loyalty and pricing show that personalized offers based on purchase history and loyalty tiers can improve gross margins by two to four percentage points compared to standard mass offers (Kumar and Shah, 2024). The critical balance involves providing sufficient value to motivate customer spending without giving away so much that costs exceed the benefits of incremental revenue (Dorotic et al., 2012).
However, implementing the comprehensive model requires measuring spillover coefficients, cannibalization rates, demand imbalances, seasonal indices, service quality scores, and customer personas. Each factor demands extensive data collection infrastructure, continuous monitoring systems, and subjective assessment frameworks that may vary substantially across brand categories, geographic markets, and competitive contexts. A complementarity discount between a coffee shop and bakery might justify a fifteen percent price reduction in one market but only five percent in another depending on local consumer preferences and competitive positioning. Similarly, the magnitude of seasonal adjustments for a hotel chain differs dramatically between leisure and business travel segments. This variability in factor calibration across contexts creates operational complexity that may obscure rather than clarify the core pricing logic.
5.2 Relaxing the Pricing Formula Based on Broker-Observable Information
To validate whether the pricing mechanism can achieve its fundamental objectives, we propose examining a relaxed version that focuses exclusively on information directly observable to the trustless broker managing the settlement layer. The broker, operating as an intermediary matching redemption requests with available liquidity, possesses complete knowledge of the order book including current asset reserves, reward balances, transaction sizes, historical exchange patterns, and pool configurations. However, the broker lacks access to external market information about brand relationships such as whether two partners compete or complement each other, demand dynamics such as seasonal patterns or temporary supply constraints, and customer characteristics such as spending history or loyalty tier status.
While the comprehensive model incorporating all eight factors would provide richer pricing intelligence when such information becomes available, our validation approach asks a more fundamental question: in the absence of these external data sources, can a pricing mechanism based solely on broker-observable information achieve the core stability and fairness objectives necessary for loyalty program success? We acknowledge that the supplementary factors identified in Section 4.2 add significant value when measurable, and in more advanced implementations these could be provided through autonomous intelligent oracles that supply verified external data to the broker. However, for this analysis we assume these oracles are dormant or unavailable, forcing us to rely exclusively on what the broker can directly observe in the order book. This relaxation serves as a validation baseline: if the mechanism succeeds with minimal information, it establishes a robust foundation upon which additional intelligence layers can be constructed. If it fails with minimal information, this reveals fundamental structural problems that additional factors might mask but not resolve.
Under the assumption that spillover, cannibalization, demand, seasonal, quality, and persona adjustments are unavailable to the broker, the comprehensive pricing model reduces to:
$P_{\text{customer}}(t)=P_{\text{base}}(t)\times\left[1+\beta_{\text{trans}}\times\max\left(0,\frac{\mu(t)-\eta}{1-\eta}\right)\right]\times\left[1+\beta_{\text{flow}}\times\phi_{A}(t)\right]$ (21)
where the transaction size factor addresses single-transaction systemic risk and the flow rebalancing factor addresses aggregate pool health. This relaxation enables us to evaluate whether these two mechanisms, drawing only on broker-observable quantities, can satisfy the validation criteria of maintaining system stability and pricing fairness.
@@ 5.3.3 Numerical Example
Research on liquidity pool management demonstrates that such threshold-based protection mechanisms, when combined with adequate initial asset provisioning, effectively prevent pool depletion while maintaining reasonable pricing for normal transaction volumes (Angeris et al., 2021). The scale invariance property ensures that small brands with modest asset reserves receive proportionally equivalent protection as large brands with deep reserves, avoiding systematic disadvantages based solely on capital availability. The broker can compute this factor using only observable quantities: the requested settlement amount $m$, the current asset balance $M_{A}(t)$, and the configured parameters $\eta$ and $\beta_{\text{trans}}$.

### v2

@@ Abstract
keywords:
Loyalty programs , mechanism design , platform economics , automated market makers , pricing mechanisms , coalition programs
journal: Electronic Commerce Research and Applications
[…]
[aff1]organization=My AI Inc.,addressline=8 The Green. Suite, city=Dover, postcode=11901, state=DE, country=USA
[…]
[aff2]organization=School of Natural and Computing Sciences, University of Aberdeen,city=Aberdeen, postcode=AB24 3FX, state=Scotland, country=UK
[…]
[aff3]organization=Marshall School of Business, University of Southern California,city=Los Angeles, postcode=90007, state=CA, country=USA
{graphicalabstract}
@@ 1 Introduction
Coalition loyalty programs emerged as the dominant attempt to implement this vision, with multiple brands partnering through central operators who manage shared currencies that customers earn and redeem across partner networks. The theoretical advantages appeared compelling, rooted in network economics, economies of scale, and enhanced customer value propositions. The empirical reality, however, proved starkly different. Industry analysis reveals that approximately 60% of coalition programs fail within 10 years of launch, with high-profile examples including American Express’s Plenti, which shut down after three years despite over \$100 million in investment (Bond Brand Loyalty, 2020; Shoulberg, 2018).
This widespread failure pattern raises a fundamental question that motivates our research. Do coalitions face merely operational challenges that better management could resolve, or do they suffer from structural problems inherent in their architectural design? Our systematic analysis reveals that coalition failures stem from fundamental architectural limitations in centralized operator models rather than operational deficiencies. These structural problems grow increasingly critical as commerce transforms, with artificial intelligence systems now mediating relationships between consumers and businesses through shopping assistants, browser extensions, and autonomous purchasing agents (Russell and Norvig, 2020). When intelligent agents require standardized, machine-readable interfaces for effective operation, the architectural limitations of traditional coalition models become absolute barriers rather than mere inefficiencies.
@@ 2.2.1 Closed Systems
However, closed systems face inherent limitations that restrict their value proposition. Customers experience fragmentation, managing multiple programs with separate accounts, balances, and rules, creating cognitive burden and reducing engagement (Nunes and Drèze, 2006). Limited network effects emerge because value remains constrained to single-brand touchpoints, preventing the accumulation synergies that multi-brand systems could theoretically provide. Smaller businesses struggle with economies of scale, finding the infrastructure investments and operational costs prohibitive relative to their customer base and transaction volumes. Breakage (unredeemed points) may be high when customers lack sufficient engagement with single brands to accumulate meaningful balances, representing dead-weight loss in the system. Most fundamentally, closed systems misapprehend the reality that users are inherently social beings. The same individual who flies an airline also dines at restaurants, shops at malls, stays in hotels, and attends sporting events. While such systems may suffice in today’s landscape, they face inevitable obsolescence as thousands of businesses compete for increasingly scarce user attention. In this fragmented environment, requiring users to independently manage rewards across disparate platforms becomes untenable. Moreover, in an emerging AI-driven agentic paradigm capable of computing complex user preferences instantaneously, closed systems dramatically underutilize optimization capabilities. Strategies that optimize across composite user goals, treating users as the social entities they are rather than isolated economic actors, will demonstrably outperform siloed approaches that fail to reflect the interconnected nature of human behavior.
@@ 2.3.1 Empirical Evidence from Coalition Failures
Air Miles Canada, one of the longest-running coalition programs, grew to over 10 million member households and more than 100 brand partners over two decades, achieving scale that should have generated the network effects and economies central to coalition value propositions. However, the program faced persistent profitability challenges that led to a critical 2016 decision where announcing that points would begin expiring triggered massive customer backlash, government scrutiny, and intense negative media coverage (CBC News, 2016a, b). The coalition operator was ultimately forced to reverse the policy, but the incident severely damaged program trust and exposed fundamental tensions between coalition economics and customer satisfaction. While Air Miles survived this crisis, the experience revealed that even large-scale, established coalitions struggle with economic sustainability.
Nectar in the United Kingdom illustrates the challenging economics underlying large coalition loyalty programs. Launched in 2002 with anchor partners Sainsbury’s, BP, Barclaycard, and Debenhams, Nectar rapidly grew to nearly 13 million members within its first year and is now described by Nectar360 as the UK’s largest coalition loyalty program (WARC, 2003, n.d.). In the years leading up to 2018, the takeover of partner Homebase by Bunnings and the exit of British Gas reduced the program’s ability to add new large non-competitive partners, and Aimia, Nectar’s operator at the time, concluded that retaining ownership offered limited opportunities to create value (Aimia, 2018b).
In February 2018, Aimia sold Nectar and related UK assets to Sainsbury’s for gross consideration of approximately Can\$105 million (£60 million), while simultaneously transferring approximately Can\$183 million (£105 million) of cash to cover redemption liabilities and approximately Can\$96 million (£55 million) of net working capital. After giving effect to the transaction, Aimia reported that its net cash and liquidity position declined by approximately Can\$174 million (Aimia, 2018b). Notably, the divested business had generated Can\$57.1 million in Adjusted EBITDA and Can\$49.9 million in Free Cash Flow during 2017 (Aimia, 2018a). The deal structure suggests that covering accumulated redemption obligations required cash transfers substantially exceeding the headline sale price. Sainsbury’s, already Nectar’s largest issuance and redemption partner across many key retail categories, was better positioned than independent operators to manage these liabilities on its own balance sheet. This transaction illustrates that even successful coalition programs can accumulate redemption obligations requiring substantial balance-sheet capacity, placing the operation of large-scale coalitions beyond the reach of firms without deep pockets.
@@
Plenti in the United States represents perhaps the most dramatic coalition failure. Backed by American Express and featuring high-profile partners including AT&T, Exxon, Macy’s, and Rite Aid, Plenti launched in 2015 with over \$100 million in marketing investment and ambitious projections to establish the first major multi-brand loyalty program in the U.S. market (Shoulberg, 2018). However, within only three years, American Express announced the program’s shutdown, citing inability to achieve projected growth and engagement targets. The rapid failure despite enormous resources, sophisticated operator expertise, and prominent partners demonstrated conclusively that coalition challenges extend beyond operational execution to fundamental structural issues that cannot be resolved through better management or increased investment.
@@ 2.5 Trustless Perpetual Systems: Architectural Foundations for Alternative Coordination
Automated Market Makers (AMMs) represent a specific smart contract application particularly relevant to reward systems. AMMs use algorithmic formulas to determine exchange rates between assets based on supply and demand reflected in liquidity pools (Adams et al., 2020). Participants provide liquidity by depositing assets, earning fees in exchange, while others execute trades at algorithmically determined rates. This mechanism enables decentralized exchange without operators setting prices, matching counterparties, or maintaining order books.
@@ 4.2.2 Dynamic Adjustment Factors
Flow Balance and Outflow Risk ($\phi_{A}$). In coalition programs, brands fund points when customers earn them. When a brand’s customers consistently redeem at partner brands without attracting corresponding return traffic, the issuing brand experiences a form of cannibalization, having invested in rewards that generate transactions elsewhere (Dorotic et al., 2021). The flow balance factor $\phi_{A}$ addresses this dynamic by providing pricing signals that discourage excessive outflow when imbalances persist, promoting network reciprocity.
Transaction Size and Market Impact ($\mu$). In liquidity-based exchange systems, large transactions relative to available pool capacity cause disproportionate price impact, affecting subsequent exchange availability and pricing stability for other participants (Angeris et al., 2021b). The transaction size factor $\mu$ addresses this dynamic by enabling pricing adjustments based on relative transaction magnitude. Larger transactions face premiums reflecting their disproportionate impact on system liquidity.
@@ 4.8.3 Numerical Example
Research on liquidity pool management demonstrates that such threshold-based protection mechanisms, when combined with adequate initial asset provisioning, effectively prevent pool depletion while maintaining reasonable pricing for normal transaction volumes (Angeris et al., 2021a). The scale invariance property ensures that small brands with modest asset reserves receive proportionally equivalent protection as large brands with deep reserves, avoiding systematic disadvantages based solely on capital availability. The broker can compute this factor using only observable quantities: the requested settlement amount $m$, the current asset balance $M_{A}(t)$, and the configured parameters $\eta$ and $\beta_{\text{trans}}$.